OpenStudy (anonymous):

FIND THE EQUATION OF A LINE GIVEN THAT: SLOPE 5/3 Y-INTERCEPT (0,-3)

OpenStudy (mef4.0):

m = 5/3 x1 = 0 y1 = -3 use point-slope form, sub in those values and you're all set point-slope is y-y1 = m(x-x1) you don't have to go any further unless you need y=mx+b form.

OpenStudy (anonymous):

i did that n i got -3=0+b but they want it in y=mx+b form

OpenStudy (mef4.0):

m = 5/3 x1 = 0 y1 = -3 use point-slope form, sub in those values and you're all set point-slope is y-y1 = m(x-x1) you don't have to go any further unless you need y=mx+b form.

OpenStudy (mef4.0):

you shud get y - (-3) = 5/3(x-0), then y + 3 = 5/3x, then subtract 3 from both sides and you're done Does that make sense?

OpenStudy (anonymous):

y=5/3 x -3

OpenStudy (anonymous):

yes somewhat thank you